OpenStudy (anonymous):

the length of a rectangle is five times its width. if the perimeter of the rectangle is 84 cm, find its area. i really have no idea how to do these...

OpenStudy (anonymous):

84=2(x+y) where x is length and y is width now 84=2(5y+y) solve it to get y then put value of y in x=5y you will have dimensions of rectangle

OpenStudy (anonymous):

im still pretty confused.. like i said i really dont get this stuff sorry..

OpenStudy (anonymous):

perimeter=2(lenght+bredth) 84=2(x+y) as given lenght = five time width so x=5y then 84=2(5y+y) puting value of x in above equation solve it you will get value of y

OpenStudy (anonymous):

ohhhh ok thank you